The data shown below represent the repair cost for a​ low-impact collision in a simple random sample of​ mini- and​ micro-vehicles. $3148 ​$1758 $1071 ​$3345 $743 $2061 ​$663 $2674 $773 $1381. Construct and interpret a 95​% confidence interval for population mean cost of repair.
Answer #1

sample mean, xbar = 1761.7
sample standard deviation, s = 1010.7736
sample size, n = 10
degrees of freedom, df = n - 1 = 9

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 2.26


ME = tc * s/sqrt(n)
ME = 2.26 * 1010.7736/sqrt(10)
ME = 722.374

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (1761.7 - 2.26 * 1010.7736/sqrt(10) , 1761.7 + 2.26 * 1010.7736/sqrt(10))
CI = (1039.33 , 2484.07)